Suppose that XYZ Corp. will generate free-cash-flows (FCF) of $300 at the end of the year. XYZ has a cost of equity capital of 15%, a cost of debt capital of 5%, a market value debt-to-equity ratio of 0.5, and faces a 21% tax rate. Assuming that XYZ’s FCF will grow by 3% per year in the future, and XYZ Corp. has in debt with a market value of $2000, what is the value of XYZ’s Corp. equity? Round your final answer to two decimals?
Answer #1

WACC = [wD * kD * (1 - t)] + [wE * kE]

= [{0.5 / (1 + 0.5)} * 5% * (1 - 0.21)] + [{1 / (1 + 0.5)} * 15%]

= 1.32% + 10% = 11.32%

Firm Value = FCF1 / (WACC - g) = 300 / (0.1132 - 0.03) = 300 / 0.0832 = $3,607.21

Value of Equity = Firm Value - Value of Debt = 3,607.21 - 2,000 = $1,607.21